OpenStudy (anonymous):

On Monday morning, Colin filled his car with six and one-half gallons of gas, then used three and one-fourth gallons driving. On Tuesday, he used up 2 gallons driving. On Tuesday night, his gas tank had four and one-fourth gallons remaining. How much gas was in Colin's tank before he filled up on Monday morning? 3 gal three and one-fourth gal three and one-half gal three and three-fourths gal

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

let x be previous fuel 6 1/2 gallons petrol filled.. 3 1/4 + 2 used fuel.. remaining gallons = 4 1/4.. so ,. x +13/4 = 2 + 9/4 + 13/2 multiply throughout by 4.. 4x + 13 = 8 + 9 + 26 4x = 30 x = 30/4 = 7 1/2

OpenStudy (anonymous):

but 7 1/2 is not one of the answers am I suppose to divide it or something @Compassionate

OpenStudy (compassionate):

What are your answers?

OpenStudy (anonymous):

3 gal , 3 1/4 gal, 3 1/2 gal and 3 3/4 gal

OpenStudy (compassionate):

Let the amount of gas in the tank before he filled up on Monday be x gallons. Activity : Amount of gas in tank Monday morning: x+6.5 Monday driving: x+6.5-3.25=x+3.25 Tuesday driving: x+3.25-2= x+1.25 = Final amount of gas. Now, use this sentence to form an equation: "On Tuesday night, his gas tank had four and one-fourth gallons remaining"

OpenStudy (anonymous):

ok so what does x =

OpenStudy (compassionate):

You're basically saying, "please, give me the answer."

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

so I just have to solve the last equation right??

OpenStudy (compassionate):

\[3\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

wow I didn't ask for the answer I asked what equation I would have to solve

OpenStudy (compassionate):

I gave you the equation and you ignored it.